OpenStudy (anonymous):

how to solve these? 7/x +4/x=1

OpenStudy (anonymous):

11

OpenStudy (austinl):

\[\frac{7}{x} + \frac{4}{x} = 1\] Is this correct?

OpenStudy (anonymous):

the same base

OpenStudy (anonymous):

yes Austin it is

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

x=4+7/1

OpenStudy (anonymous):

1/11=11=x

OpenStudy (austinl):

Rewind, He is correct. \[\frac{7+4}{x} = 1\rightarrow\frac{7+4}{1}=x=11\]

OpenStudy (austinl):

For the record, @yahya90 there are more polite ways to point out that someone is incorrect. That was rude.

OpenStudy (anonymous):

I agree Austin- sorry yahya-that was rude

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i am realy sorry

OpenStudy (anonymous):

\[\frac{ 7 }{ x }+\frac{ 4 }{ x }=1\] \[\frac{ 7+4 }{ x }=1\] \[7+4 = x\] \[11 = x\] so x = 11 :)

OpenStudy (anonymous):

thank you guys

OpenStudy (austinl):

No problem!